Find the expected count and the contribution to the chi-square statistic for the (C,E) cell in the two-way table below

D E F G Total

A 35 30 44 41 150

B 78 90 67 58 293

C 19 38 26 32 115

Total 132 158 137 131 558

Round your answer for the expected count to one decimal place, and your answer for the contribution to the chi-square statistic to three decimal places.

Expected count =

contribution to the chi-square statistic =
Answer #1

1)since expected count =row total*column total/grand total=(115*132/558)=27.2

2) contribution to the chi-square statistic = (O-E)2/E =(19-27.2)^2/27.2=2.472